In a class, 70 percent are male. Five new female students enrolled in the class and the percentage of the male student become 60 percent. What was the total original number of students in the class? Use x to represent the total number of original students.
Answer by sachi(548)

In a class, 70 percent are male. Five new female students enrolled in the class and the percentage of the male student become 60 percent. What was the total original number of students in the class? Use x to represent the total number of original students.

ans:-
let there are x students in toto
male students=0.7x
female students=0.3x
5 female students joined
so the total students=x+5
the % of male students=(0.7x)/(x+5) =0.6 as per question
or 7x=6(x+5)=6x+30
or x=30 the total no of students ans